OpenStudy (anonymous):

PLEASE HELP. WILL MEDAL ~ Solve. 3y + 15 > –6 A. y > 3 B. y > –7 C. y > 13 D. y > –17

OpenStudy (dmb1215):

Ok first you need to -15 from both sides

OpenStudy (dmb1215):

so that would give you 9

OpenStudy (dmb1215):

then what is 9/3

OpenStudy (anonymous):

9 divided by 3?

OpenStudy (dmb1215):

yep

OpenStudy (anonymous):

3

OpenStudy (anonymous):

so your answer is A

OpenStudy (dmb1215):

there is you answer

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

np

OpenStudy (dmb1215):

np

OpenStudy (anonymous):

wouldn't it be? 3y +15 > -6 -15 -15 3y > -21 -------- 3 3 y > -7

OpenStudy (dmb1215):

what is a negative - a negative

OpenStudy (dmb1215):

if your subtracting from a negative it means your adding

OpenStudy (anonymous):

i keep getting -21

OpenStudy (anonymous):

i think you're thinking of multiplication rule for -1 time -1 = 1

OpenStudy (anonymous):

subtraction Ex. 6 - 2 = 6 + (-2) = 4 Ex. -6 - -2 = -6 + (+2) = -4 Ex. -6 - 2 = -6 + (-2) = -8 <----- Ex. 6 - -2 = 6 + (+2) = 8 so this would make the problem -21 since -6 - 15 forgive me if im wrong, its been years since i did this

OpenStudy (anonymous):

i did a little research http://mathhomeworkanswers.org/49400/what-is-y-in-3y-15-6

OpenStudy (triciaal):

Rule of thumb always picture the number line when you add go to the right when you subtract go to the left when you add a minus you are subtracting when you subtract a minus you are "bouncing back" so add go right for this problem 3y + 15 > –6 what @HeatDude has is correct ; y>-7 to check you can always pick a number and plug it in example use a number >-7 choose 2 then 3*2 + 15 = 6 + 15 =21 which is >-6 example use a number <-7 choose -8 then 3*(-8) +15=-24 + 15 which is <-6

OpenStudy (dmb1215):

yes i was wrong so go ahead and beat me down im sorry